# OpenECAD: An Efficient Visual Language Model for Computer-Aided Design

### v2

@@ 4.6 Creation of CAD Dataset
Based on the above method, we generated a 3D CAD dataset derived from DeepCAD dataset[24] and Fusion 360 Gallery Reconstruction Dataset[26]. We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the number of tokens that small language models can accept is limited, resulting in a total of 130,239 3D designs. The statistical distribution of the number of ”Sketch-Extrusion” steps in the 3D dataset is shown in the Figure 5.
Figure 5: The statistical distribution of the number of ”Sketch-Extrusion” steps in OpenECAD 3D dataset.
@@ 5.1 The choice of language models and multimodal approach
Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[28] (450M), Gemma[29] (2B), Phi-2 (2.7B) and Phi-3 mini[30] (3.8B) with relatively small parameter sizes as the foundation for our language model. OpenELM is an efficient language model family developed by Apple. It employs a layer-wise scaling strategy to allocate parameters efficiently within each layer of the transformer model. This approach enhances accuracy while maintaining computational efficiency. Gemma is a family of lightweight, text-to-text, decoder-only large language models developed by Google. Phi, developed by Microsoft Research, is another series of advanced small language models (SLMs). This series includes Phi-2 (2.7B), Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B), each with different parameter sizes. Despite their compact parameter scales, Phi-3 models achieve language understanding and reasoning capabilities comparable to larger models. These models are trained on high-quality data and optimized algorithms. It is worth noting that the context acceptance length for OpenELM (450M) is 2048 tokens. In contrast, Gemmma and Phi has a minimum context acceptance length of 3072 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
To endow the language model foundation with multimodal capabilities, we selected OpenAI’s CLIP[34] and Google’s SigLIP [35] as visual encoders for OpenELM (450M), Gemma (2B) and Phi-2 (2.7B). These models are trained on large datasets consisting of images paired with corresponding text descriptions, combining visual and textual information to make predictions.
To train the entire model with multimodal conversational abilities, we employed the LLaVA[36] method and its dataset. LLaVA combines a visual encoder and a large language model in an innovative multimodal approach, enabling comprehensive visual and language understanding. LLaVA uses GPT-4 to generate multimodal language-image instruction data. For training the model, we utilized the training framework provided by TinyLLaVA[31].
For Phi-3 mini, since Microsoft has already released its multimodal version, Phi-3 Vision (4.2B), which was trained using methods similar to LLaVA, we will directly use the pre-trained model provided by Microsoft.
@@ 5.2 Fine-tuning on a pre-trained visual language model
OpenECAD 0.55B & 0.89B
CPU Intel Core i9-7960X
GPU Nvidia GeForce GTX 1080 Ti
RAM DDR4-2666 32 GB
@@
OpenECAD 2.4B & 3.1B
CPU AMD EPYC 9754 Processor (18 vCPU)
GPU 2 x Nvidia GeForce RTX 4090D
RAM 60 GB
OS Ubuntu Server 22.04 LTS
OpenECAD 4.2B
CPU AMD EPYC 9754 Processor (18 vCPU)
GPU Nvidia GeForce RTX 4090D
RAM 60 GB
@@
We performed LoRA fine-tuning on the pre-trained OpenELM-CLIP (0.55B) and OpenELM-SigLIP (0.89B) models using an Nvidia GeForce GTX 1080 Ti (11 GB VRAM). The Rank and Alpha for LoRA were set to 128 and 256, respectively. The learning rate is $10^{-4}$. Training for one epoch with our 3D dataset took approximately 60 hours. Through fine-tuning, we obtained the OpenECAD 0.55B and 0.89B models.
Similarly, we used 2 Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Gemma-SigLIP (2.4B) and Phi-2-SigLIP (3.1B) model. The Rank and Alpha for LoRA were set to 128 and 256, respectively. The learning rate is $10^{-4}$. Training for one epoch with our 3D dataset took approximately 14 hours. Through fine-tuning, we obtained the OpenECAD 2.4B and 3.1B models.
We used an Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Phi-3 Vision (4.2B) model. The LoRA rank and alpha were set to 256 and 512, respectively. The learning rate is $10^{-4}$. Training with our dataset for one epoch took approximately 20 hours. Through fine-tuning, we obtained OpenECAD 4.2B model.
The specific development, training, and runtime environments are shown in the Table 2. The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training are illustrated in the Figure 6. Due to the different batch sizes used during training, the loss curves are proportionally scaled. As shown, the models converge after a single round of fine-tuning.
Figure 6: The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training.
@@ 6 Experiments
In this chapter, we evaluate the ability of OpenECAD models to generate CAD designs based on input images and analyze the results. As there are no existing CAD models that use images for CAD generation, we first define evaluation metrics (Section 6.1) and create CAD designs for testing (Section 6.2). Following this, we test the models, obtain the results (Section 6.3), and analyze them (Section 6.4).
@@ 6.1 Evaluation Metrics
Is the code executable? 10
Accuracy of 3D shape construction 90
For each Loop-Extrusion pair (out of 100)
- 2D sketch positioning 20
- Accuracy of 2D Loop 60
- 3D operation 20
@@ 6.1.1 Scoring Algorithm for Evaluating Outputs
To assess the generative capability of our model, we designed a scoring algorithm. First, the algorithm verifies if the generated code can be executed directly by the API operator to create a model. If it cannot run, it scores 0 points. If it can run, we then use a visual language model (such as GPT-4o) to compare the generated model’s view with the target input, evaluating each basic feature. Each basic feature corresponds to a 2D loop. We separately evaluate the loops in the 2D sketches during the modeling process to ensure precise assessment, as these loops may or may not be considered a single sketch during modeling.
We score the results based on executability and the correctness of the 3D shapes (see Table 3). Each Loop-Extrusion pair is considered a basic 3D shape. If the number of generated Loop-Extrusion pairs is insufficient, points will be deducted proportionally. For each Loop-Extrusion pair, we score based on the reference plane position of the 2D sketch, the extrusion operation, and the drawing of the 2D loop. The specific calculation method is shown in the following formula.
$score=10R+\frac{90}{N_{s}}\sum_{i=1}^{N_{s}}(20P_{i}+20O_{i}+60L_{j})$ (1)
where $R$ denotes executability, $N_{s}$ represents the number of Loop-Extrusion pairs, $P_{i}$ and $O_{i}$ indicate the correctness of the 2D sketch position and 3D extrusion operation for the $i$-th Sketch-Extrusion pair, and $L_{j}$ denotes the accuracy of the curves within the loop. $R,P_{i},O_{i},L_{j}\in[0,1]$.
@@ 6.2 CAD Designs for Test
To test the model, we require high-quality CAD designs. We extracted approximately 100 3D designs from the training set, maintaining a ratio of 10:5:3:2 for designs containing 1, 2, 3, and 4 or more Sketch-Extrusion pairs, respectively. After filtering and removing duplicates, we obtained 57 unique 3D designs. These designs were redrawn using CAD tools such as AutoDesk Fusion, and appropriate views were selected to generate the input images. Some examples are shown in the Figure 7. The resolution of the images is 640x400.
@@ 6.3 Evaluation Results
Figure 8: Comparison of some successful rendered test results illustrations.
@@ 6.3.1 The Inference Speed of the Models
The input was standardized to a 640x400 image of a 3D shape, with the prompt:
”This image is a view of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
For OpenECAD 0.55B and 0.89B, we used an Nvidia GeForce GTX 1080 Ti (11 GB VRAM) for inference, with a peak memory usage of less than 5GB. Getting a 3D shape’s code took approximately 20 seconds. For OpenECAD 2.4B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 9GB. For OpenECAD 3.2B and 4.2B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 18GB. Getting a 3D shape’s code took approximately 45 seconds.
@@ 6.3.2 Generating Code and Rendering
We tested OpenECAD versions 0.55B, 0.89B, 2.4B, 3.2B and 4.2B, and also tested GPT-4o. Some successfully rendered test results are illustrated in Figure 8. Since GPT-4o was not trained specifically for the use case described in this paper, we used the following prompt for testing:
”I am going to provide you with a 3D shape’s corresponding Python code for its modeling process. After that, I will input images of other 3D shapes, and please provide the corresponding Python code for them. $<$OpenECAD style Code Example$>$ This image is a view of a 3D model from a certain angle. Please write the codes of it in the format above.”
During the testing of OpenECAD, the maximum number of new tokens that 0.55B, 0.89B, 3.2B and 4.2B models can generate is 1536, 1024, 2048, 3072 and 3072, respectively. Due to these context limitations, there are instances where the generated code is incomplete. GPT-4o, due to its support for a longer context, did not encounter issues with incomplete code generation.
@@
Based on the scoring method mentioned above, we obtained the scores, the number of executable results, the number of incomplete results, the number of runtime errors, and the number of completely correct results for each model, as shown in the Table 4.
Model Overall Executable Unfinished Run with Completely
Score Errors Correct
0.55B 46.52 36 21 0 15
0.89B 37.19 22 35 0 19
2.4B 59.86 50 7 0 20
3.1B 55.39 53 0 4 16
4.2B 51.91 51 4 2 14
GPT-4o 15.44 52 0 5 4
@@
As shown in Table 4, OpenECAD 2.4B achieved the highest overall score and the most fully correct outputs. Both 0.55B and 0.89B models produced a significant amount of incomplete code. The 0.89B model’s higher token consumption for image inputs reduced the number of new tokens it could generate, leading to more incomplete outputs compared to the 0.55B model, and thus a lower overall score. The 3.1B and 4.2B models achieved relatively high scores, but the number of fully correct generations was lower than that of the 0.55B and 0.89B models. All OpenECAD models demonstrated significant advantages over the unrefined GPT-4o, whose sole advantage was its ability to handle the entire code due to its extended context length.
@@ 6.4.2 Generation Error Analysis
Figure 9: Typical errors in generated results.
In the generated results, several typical errors were observed (see Figure 9), including:
  • 1.
    Case 1 and Case 2: For complex and irregular shapes, the corresponding sketches often fail to generate correctly, or the profiles of the 2D sketches are inaccurately created. Possible reasons include the vision tower’s low resolution, which hinders accurate recognition of complex shapes, and the language model’s inability to comprehend complex shapes from the CAD operation code. These issues ultimately result in incorrect loops or intersecting loops, leading to erroneous sketch profiles and broken extrusions.
  • 2.
    Case 3 and Case 7: The model lacks spatial reasoning capabilities, resulting in incorrect handling of the size proportions between sketches and extrusions. This leads to extrusions that are either too long or too small, causing discrepancies in size relative to other parts.
  • 3.
    Case 4: When multiple ”Sketch-Extrusion” steps are involved, the positions of the sketches in subsequent steps are often incorrect, leading to errors in the final overall model.
  • 4.
    Case 5: Due to the minimal color difference between the outer surface of a solid cylinder and the inner wall of a hollow cylinder, the primary distinguishing feature is the inner wall circle at the top. This makes it difficult for the model to differentiate between solid and hollow cylinders, leading to confusion. This issue also extends to other similar shapes with holes.
  • 5.
    Case 6: The difficulty in distinguishing concave and convex shapes results in erroneous models. This issue may arise because the vision tower can detect these shapes but cannot differentiate between them, or because the language model fails to generate the correct code, leading to sketch errors.
  • 6.
    Case 8: Missing part(s). For some relatively subtle features, the model struggles to accurately identify them, resulting in missing features.
Based on the above results and typical errors, we can infer the current shortcomings of the OpenECAD model:
  • 1.
@@
  • 3.
    For complex shapes involving multiple ”Sketch-Extrusion” steps, the model struggles to generate accurate results. It has difficulty remembering the previously drawn parts through the context, leading to errors.
  • 4.

### v3

@@ 4.6 Creation of CAD Dataset
Based on the above method, we generated 3 OpenECAD datasets derived from the DeepCAD dataset. They contain 100,000, 150,000, and 200,000 “image-code” pairs, respectively, with code lengths limited to 1024 tokens, 2048 tokens, and 3072 tokens. We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the limited number of tokens that small language models can accept, as well as designs used for the validation set.
The distribution of “Default View”, “Transparent View”, and “Orthographic Views” in the datasets is shown in the Table 2. “$x$_directout_$y$” indicates that the dataset contains $x$ “image-code” pairs, with the code length limited to $y$ tokens. This dataset is used for directly outputting complete OpenECAD code from image input.
Dataset Name Default View Transparent View Orthographic Views
100k_directout_1k 62500 25000 12500
150k_directout_2k 90000 37500 22500
200k_directout_3k 120000 50000 30000
Table 2: The distribution of OpenECAD datasets.
The statistical distribution of the number of “Sketch-Extrusion” steps in the OpenECAD datasets is shown in the Figure 4. As shown in the figure, designs with only one “sketch-extrude” pair still constitute the majority of datasets, especially in datasets with strict token limits. This is partly because the token limit prevents OpenECAD code from being long enough to accommodate more “sketch-extrude” pairs in a single design, and partly because such designs make up the majority of the DeepCAD dataset. In datasets where the token limit is less restrictive, more complex designs have a more significant proportion.
Figure 4: The statistical distribution of the number of “Sketch-Extrusion” steps in OpenECAD datasets.
@@ 5.1 The choice of language models and multimodal approach
Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[26] (450M), Gemma[27] (2B) and Phi-2 (2.7B)[28] with relatively small parameter sizes as the foundation for our language model. These models are trained on high-quality data and optimized algorithms. It is worth noting that the context acceptance length for OpenELM (450M) is 2048 tokens. In contrast, Gemma and Phi has a minimum context acceptance length of 3072 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
To endow the language model foundation with multimodal capabilities, OpenAI’s CLIP[24] and Google’s SigLIP are selected [25] as visual encoders for OpenELM (450M), Gemma (2B) and Phi-2 (2.7B). These models are trained on large datasets consisting of images paired with corresponding text descriptions, combining visual and textual information to make predictions.
To train the entire model with multimodal conversational abilities, the LLaVA[30] method and its dataset are employed. LLaVA combines a visual encoder and a large language model in an innovative multimodal approach, enabling comprehensive visual and language understanding. LLaVA uses GPT-4 to generate multimodal language-image instruction data. We utilized the training framework and the pretrained models provided by TinyLLaVA[29].
@@ 5.2 Fine-tuning on a pre-trained visual language model
OpenECAD 0.55B & 0.89B & 3.1B
CPU Intel Xeon Platinum 8352V
GPU 2 x Nvidia GeForce RTX 4090
RAM 90 GB
@@
OpenECAD 2.4B
CPU Intel Xeon Platinum 8352V
GPU 2 x Nvidia GeForce RTX 4080 Super (as one vGPU)
RAM 90 GB
@@
The specific development, training, and runtime environments are shown in the Table 3. The loss function used was CrossEntropyLoss, and the training method was instruction fine-tuning. The learning rate was set to $10^{-4}$, and the learning rate scheduler type was cosine. The training of all OpenECAD models were conducted with a train batch size of 2 per device and gradient accumulation steps of 2. The Rank and Alpha for LoRA were set to 128 and 256. Besides using LLM to add comments to a small portion of OpenECAD datasets’ codes for data augmentation, further data augmentation is of limited significance since the datasets are already sufficiently large for fine-tuning.
We performed LoRA fine-tuning on the pre-trained OpenELM-CLIP (0.55B), OpenELM-SigLIP (0.89B) and Phi-2-SigLIP (3.1B) models. The training was conducted on 2 Nvidia GeForce RTX 4090 GPUs. Under the environment described in Table 3, fine-tuning on the pre-trained OpenELM-CLIP (0.55B) and OpenELM-SigLIP (0.89B) models with the OpenECAD 100k_directout_1k dataset took approximately 2 hours, fine-tuning on the pre-trained Phi-2-SigLIP (3.1B) model with the OpenECAD 200k_directout_3k dataset took approximately 20 hours.
Similarly, we used 2 Nvidia GeForce RTX 4080 Super GPUs as a single vGPU to perform LoRA fine-tuning on the Gemma-SigLIP (2.4B) model. Under the environment described in Table 3, fine-tuning on the pre-trained Gemma-SigLIP (2.4B) model with the OpenECAD 150k_directout_2k dataset took approximately 30 hours
The loss curves of the first 1500 steps for the OpenECAD 0.55B, 0.89B, 2.4B, and 3.1B models during training are illustrated in Figure 5. As shown, the models converge after the first 1500 steps of fine-tuning.
Figure 5: The partial loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, and 3.1B models during training.
@@ 6 Experiments
In this chapter, we evaluate the ability of OpenECAD models to generate CAD designs based on input images and analyze the results. As there are no existing CAD models that use images for CAD generation, we first define evaluation metrics (Section 6.1) and create CAD designs for testing (Section 6.2). Following this, we test the models, obtain the results (Section 6.3), and analyze them (Section 6.4). We also provide an example of drawing a simple table (Section 6.5) and discuss future work (Section 6.6).
@@ 6.1 Evaluation Metrics
Is the code executable? 10 * exec_flag
Accuracy of All Curves 45 * acc
Accuracy of Loops count 5 * acc
Loops integrated score 40 * score
Fo each Loop’s score (out of 100)
If loop is absolutely correct 100
Else calculate acc of curves 90 * acc
@@ 6.1.1 Scoring Algorithm for Evaluating Outputs
To assess the generative capability of our model, we designed a scoring algorithm. The algorithm verifies if the generated code can be executed directly by the API operator to create a model, the accuracy of curve operations, the accuracy of the number of loops, and provides a comprehensive score for the loops. We separately evaluate the loops in the 2D sketches during the modeling process to ensure precise assessment, as these loops may or may not be considered a single sketch during modeling. We score the results based on executability and the correctness of the 3D shapes (see Table 4).
For the comprehensive scoring of loops, we first check if the two loops are completely identical, meaning the types and order of the curves are exactly the same, and assign a score of 100. If not, the score is recorded as 90% of the curve accuracy. The specific calculation method is shown in the following formula.
$score=10e+45acc_{c}+5acc_{l}+\frac{40}{100L}\sum_{i=1}^{L}(10s_{i}+90acc_{i})$ (1)
where $e$ denotes executability, $acc_{c}$ represents accuracy of curve operations, $acc_{l}$ represents accuracy of the number of loops, $L$ represents the number of Loops, $s_{i}$ and $acc_{i}$ indicate whether the loop is absolutely correct and the accuracy of curve operations for the $i$-th loop. $e,acc_{c},acc_{l},s_{i},acc_{i}\in[0,1]$.
@@ 6.2 CAD Designs for Test
To test the model, we require high-quality CAD designs. We extracted 625 3D designs from the OpenECAD dataset that were not included in the training set. Similar to the method used for generating the training set, we created their standard views, perspective views, and four-angle views. Some examples are shown in the Figure 6. The resolution of the images is 640x480. The statistical distribution of the number of “Sketch-Extrusion” steps in the is shown in the Figure 7.
Figure 7: The statistical distribution of the number of “Sketch-Extrusion” steps of CAD designs for evaluating.
@@ 6.3.1 The Inference Speed of the Models
The input was standardized to a 640x480 image of a 3D shape, with the prompt:
“This image is a(4) view(s) of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
For all OpenECAD models, we used an Nvidia GeForce GTX 1080 Ti (11 GB VRAM) for inference. OpenECAD 0.55B and 0.89B models cost a peak memory usage of less than 4GB. Getting a 3D shape’s code took approximately 30 seconds. For OpenECAD 2.4B, it infers with a peak memory usage of less than 9GB. OpenECAD 3.1B infers with a peak memory usage of less than 11GB. Getting a 3D shape’s code took approximately 45 seconds using OpenECAD 2.4B and 3.1B models.
@@ 6.3.2 Generating Code and Rendering
Figure 8: Some successful rendered test results illustrations
We tested OpenECAD versions 0.55B, 0.89B, 2.4B and 3.1B, and also tested GPT-4o-mini (2024-07-18 version). Some successfully rendered test results are illustrated in Figure 8. Since GPT-4o-mini was not trained specifically for the use case described in this paper, we used the following prompt for testing:
“I am going to provide you with a 3D shape’s corresponding Python code for its modeling process. After that, I will input images of other 3D shapes, and please provide the corresponding Python code for them. $<$OpenECAD style Code Example$>$ This image is a view of a 3D model from a certain angle. Please write the codes of it in the format above.”
During the testing of OpenECAD, the maximum number of new tokens that the 0.55B, 0.89B, 2.4B, and 3.1B models can generate is 1536, 1152, 2048, and 2048 respectively. Due to these context limitations, there are instances where the generated code is incomplete. GPT-4o-mini, due to its support for a longer context, did not encounter issues with incomplete code generation.
@@
Based on the scoring method mentioned above, we obtained Curves accuracy, Loops count accuracy, Loops score, Execution score, and Overall score for each model and input type, as shown in the Table 5. Additionally, we calculated the scores excluding designs with only one “Sketch-Extrusion” pair.
Default View as input Default View without only 1 pair designs
Model 0.55B 0.89B 2.4B 3.1B gpt-4o-mini 0.55B 0.89B 2.4B 3.1B gpt-4o-mini
Curves accuracy 85.95 88.97 91.23 87.71 40.11 68.65 74.27 76.01 70.29 25.81
Loops accuracy 84.27 86.79 90.62 86.98 75.99 74.23 76.72 76.43 70.63 57.10
Loops score 77.73 80.03 84.20 81.46 34.45 65.50 68.57 72.96 69.18 25.07
Execution score 99.04 96.64 95.36 94.40 40.96 98.92 90.32 85.48 84.95 24.73
Overall score 83.89 86.05 88.80 85.84 39.72 70.70 73.72 75.76 71.33 26.97
Transparent View as input Transparent View without only 1 pair designs
Model 0.55B 0.89B 2.4B 3.1B gpt-4o-mini 0.55B 0.89B 2.4B 3.1B gpt-4o-mini
Curves accuracy 84.57 88.03 90.68 86.72 47.38 70.00 73.65 78.52 70.78 30.00
Loops accuracy 83.12 84.98 88.74 85.90 76.20 75.80 76.49 76.57 73.36 58.54
Loops score 77.57 79.83 83.94 81.32 41.02 66.68 68.71 75.41 71.71 28.88
Execution score 99.20 97.76 95.04 93.44 44.48 99.46 93.55 85.48 82.26 25.81
Overall score 83.16 85.57 88.32 85.19 45.99 71.91 73.81 77.88 72.43 30.56
Orthographic View as input Orthographic View without only 1 pair designs
Model 0.55B 0.89B 2.4B 3.1B gpt-4o-mini 0.55B 0.89B 2.4B 3.1B gpt-4o-mini
Curves accuracy 85.28 87.02 90.34 87.71 31.78 67.63 72.08 74.91 69.32 22.96
Loops accuracy 82.73 84.46 89.17 87.02 71.11 70.31 75.17 73.89 70.08 54.86
Loops score 76.50 79.02 83.21 81.14 26.75 63.18 68.47 72.15 68.50 21.97
Execution score 99.68 97.44 95.52 95.20 34.88 98.92 91.94 86.56 84.41 23.12
Overall score 83.08 84.73 87.95 85.80 32.04 69.11 72.78 74.92 70.54 24.17
@@
As shown in Table 5, OpenECAD 2.4B achieved the highest overall score. The 0.55B and 0.89B models, despite generating the fewest tokens, scored relatively high in executability. This might be because their inability to understand complex models leads them to finish code generation prematurely. The 3.1B model did not perform as expected, possibly due to the Phi series models having inferior geometric understanding compared to the Gemma series models. All OpenECAD models demonstrated significant advantages over the unrefined GPT-4o-mini, whose sole advantage was its ability to handle the entire code due to its extended context length.
From the perspective of using different types of view as input, using single-view transparent images as input shows a significant advantage in designs with more than one “sketch extrusion” pair because they contain more information. However, using four views as input (isometric view + three orthographic views) performs worse, possibly because the vision tower cannot handle so much information, or the language model cannot establish connections between them. The additional views increase the amount of irrelevant information, adding pressure to the model.
@@ 6.4.2 Generation Error Analysis
In the generated results, several typical errors were observed in the Table 6.
Error Case (Left: target, Right: generation result) Analysis
[Uncaptioned image] For complex and irregular shapes, the corresponding sketches often fail to generate correctly, or the profiles of the 2D sketches are inaccurately created. Possible reasons include the vision tower’s low resolution, which hinders accurate recognition of complex shapes, and the language model’s inability to comprehend complex shapes from the CAD operation code. These issues ultimately result in incorrect loops or intersecting loops, leading to erroneous sketch profiles and broken extrusions.
[Uncaptioned image] The model lacks spatial reasoning capabilities, resulting in incorrect handling of the size proportions between sketches and extrusions. This leads to extrusions that are either too long or too small, or sketches that are either too large or too small, causing discrepancies in size relative to other parts. The lack of numerical understanding is a common issue with language models, which can be relatively easily resolved by post-editing the extent size of the extrusions or the size of curves in the sketches in the CAD code or project.
[Uncaptioned image] When multiple “Sketch-Extrusion” steps are involved, the positions of the sketches in subsequent steps are often incorrect, leading to errors in the final overall model. This can be resolved by post-editing the reference or absolute positions of the sketches in the CAD code or project. However, for complex models, this solution might be more challenging.
[Uncaptioned image] Due to the minimal color difference between the outer surface of a solid cylinder and the inner wall of a hollow cylinder, the primary distinguishing feature is the inner wall circle at the top. This makes it difficult for the model to differentiate between solid and hollow cylinders, leading to confusion. This issue also extends to other similar shapes with holes.
[Uncaptioned image] The difficulty in distinguishing concave and convex shapes results in erroneous models. This issue may arise because the vision tower can detect these shapes but cannot differentiate between them, or because the language model fails to generate the correct code, leading to sketch errors.
Table 6: Error Cases and Related Analysis
Based on the results and typical errors mentioned in the Table 6, we can infer the current shortcomings of the OpenECAD model:
@@
  • For complex shapes involving multiple “Sketch-Extrusion” steps, the model struggles to generate accurate results. It has difficulty remembering the previously drawn parts through the context, leading to errors.
@@
6.5 An Example of Using OpenECAD models
Figure 9: An example of using OpenECAD 2.4B model to draw a table.
In this section, we will attempt to use the OpenECAD model by inputting a hand-drawn simple table (see Figure 9), generating the corresponding OpenECAD code, changing the length of 4 legs of the table and then converting it into a step file using the PythonOCC API.
We input a hand-drawn table’s photo to the OpenECAD 2.4B model, with the prompt:
“This image is a hand-drawn view of a 3D model from a certain angle. Please try to use Python-style APIs to render this model.”
The model should respons:
“Of course, here are the codes:$<$OpenECAD code here$>$
Algorithm 3 Some OpenECAD Code of the Table
1:  SketchPlane0 = add_sketchplane(origin= [0., 0., 0.], normal= [0., 0., 1.], x_axis= [1., 0., 0.]))
2:  Loops0, Curves0_0 = [], []
3:  Line0_0_0 = add_line(start= [-1000., -750.], end= [1000., -750.])
4:  …
5:  SketchPlane1 = add_sketchplane_ref( Extrude0, origin = [0., 0.], type = “sameplan”, reverse = True)
6:  …
7:  Extrude1 = add_extrude(sketch= Sketch1, operation= 1, type= 0, extent_one= 1000., extent_two= 0.)
8:  {extent_one is the length of one of 4 legs of the table.}
Some key code snippets are shown in the Algorithm 3. We changed the extent_one of the table legs extrusion from 1000 to 500. We converted the code before and after the change into STEP files using a conversion tool and opened them with a CAD tool. As shown in the Figure 9, the table legs are indeed shortened by half.
Additionally, besides directly modifying the OpenECAD code, it is also possible to convert the OpenECAD code into project files for other 3D CAD tools. For instance, AutoDesk Fusion supports all operations in OpenECAD. We can use tools and the AutoDesk Fusion API to convert OpenECAD code into AutoDesk Fusion project files. This way, we can fully utilize all the features of AutoDesk Fusion to continue modifying the design. Due to the limitations of the model’s context length, it is currently challenging for the model to directly modify the generated OpenECAD code based on requirements through dialogue, as the context length does not support accommodating two complete OpenECAD code segments.